Let k represent the length of a rectangle. Represent the perimeter when the width is 8 less than the length.
(A) 6k-6
(B) 2k+16
(C) 4k-10
(D) 4k-16
Also, can you try to tell me how you got that answer. thank you so much! I am stuck on this problem for a long time and I just found out about this website. Hopefully you can answer this question

Answer by lwsshak3(11628)

Let k represent the length of a rectangle. Represent the perimeter when the width is 8 less than the length.
***
k=length
k-8=width
perimeter=2(length+width)=2(k+k-8)=4k-16 (ans D)